# Attention Pruning: Automated Fairness Repair of Language Models via Surrogate Simulated Annealing

### v2

@@ 4. RQ4
Model Baseline AP ($\epsilon=0.5$) FASP (Zayed et al., 2024) ($\gamma=0.3$) Weight Magnitude (Han et al., 2016) Gradient Magnitude (Michel et al., 2019) Random
@@
Table 6 presents the bias and perplexity of the baseline LLM, AP, and the other strategies. The values in bold correspond to the best results. We see that in 4/6 cases, AP finds states with better bias and perplexity than FASP. The largest absolute and relative improvement in the bias of $0.18$ and $40.8\%$ is observed for GPT-J-6B. For GPT-Neo-125M, AP’s bias is worse by $\approx 0.02$, but the perplexity is better by $\approx 6$. For Llama-2-7B, AP’s bias is better by $\approx 0.03$ and the perplexity is worse by $\approx 0.3$. Overall, in most cases, AP can reduce bias more while ensuring better perplexity. The other strategies, which do not account for bias, are worse than both AP and FASP [^3] Gradient Magnitude for Llama-2-7B is left blank as its implementation in Huggingface does not provide an option for computing the gradient of masked heads..
@@
  • The cost function in SA is a linear weighted combination of the two surrogate DNNs. The hyperparameter $\epsilon$ determines how much weight to give to each DNN. Higher values of $\epsilon$ assign more weight to the bias DNN, prioritizing the reduction of bias over the loss of perplexity during the SA search. To visualize the bias-perplexity tradeoff in LLMs, we vary $\epsilon\in[0.3,0.7]$ during the SA search for four LLMs. Figure 2 shows plots of bias versus perplexity of the states for different epsilon values. States with low bias and high perplexity occur in the upper left region of the graph. States with high bias and low perplexity occur in the lower right region. From the plots, except for outliers that occur due to the randomness of SA, we see that the smaller bubbles that represent lower values of $\epsilon$ occur in the lower right region. Similarly, the larger bubbles moving towards the upper right region represent larger values of $\epsilon$.
@@
  • During the creation of the dataset for training the surrogate DNNs, increasing $\eta_{bias}$ and $\eta_{ppl}$ will use larger portions of the HolisticBias and WikiText-2 datasets. This can lead to more accurate DNN predictions as larger subsets tend to accurately represent the true bias/perplexity of the entire dataset. To analyze how accurately the subsets represent the true bias, we sample 25 subsets each for different values of $\eta_{bias}\in[0.1,0.9]$ and plot the error in bias. Figure 3 shows a plot of the error versus $\eta_{bias}$. The dotted line is a regression line fitting all points. Overall, we observe a positive trend where the error decreases as $\eta_{bias}$ is increased i.e., larger portions of the validation dataset are used to compute the bias of each sampled subset.
@@
  • Increasing the running time (time_limittime\_limit) of the SA search may lead to better states as the search can explore more states. To analyze the effect of running time on the cost and bias/perplexity of best states found, we run the SA search for 5.5 hours with $\epsilon=0.5$ and record the best result at 30-minute intervals. Figure 4 highlights the results for GPT-J-6B. From Figure 4(a), we see that the cost of the best state is a decreasing function. The cost is guaranteed to remain the same or decrease over time as the SA search explores more regions of the search space. However, bias/perplexity has a more complicated relationship. From Figure 4(b), we see that bias and perplexity alternate over time, eventually ending with the bias improving more at the expense of perplexity. We observe this because the cost function is a weighted linear function of bias and perplexity. From Figure 2, we see that the best option to control the bias/perplexity trade-off is to vary $\epsilon$.
@@
Answer RQ3: Using larger portions of the bias and perplexity datasets leads to more accurate estimates of bias/perplexity that can improve the DNN performance. Increasing the weightage given to the bias DNN in the cost function reduces bias at the expense of perplexity. Running SA longer may lead to better solutions.
@@
One of the possible unintended side effects of mitigating one form of bias is that it may exacerbate other forms of bias in the LLM. The HolisticBias metric contains prompts for 13 social biases, and we primarily focus on mitigating gender bias. Similar to Zayed et al. (2024), to understand how the other social biases are affected by AP, we pick 4 other social biases and evaluate them after repairing gender bias. We pick the best subset of attention heads AP found for gender bias, prune the LLMs, and finally evaluate the repaired LLM on the other biases. Table 7 highlights the effect on other biases before and after repairing gender bias. Note that the perplexity remains identical to the values reported in Table 6 as the perplexity is always evaluated with the WikiText-2 test dataset. We observe that all other social biases decrease for all LLMs. We see a consistent improvement in all cases, even for biases such as age, which are low by default. The highest absolute improvement in bias of 0.29 is observed for GPT-Neo-125M for sexual orientation bias. The highest relative improvement of $65.21\%$ is observed for GPT-J-6B in the age bias category.
@@ 6. Related Works
The SE community has extensively focused on mitigating unfairness in DNNs, and the literature on mitigating unfairness in LLMs is sparse. DICE (Monjezi et al., 2023) uses a fault localization technique via do logic to identify the top $k$ neurons negatively affecting fairness. FASP uses a similar approach to analyze the effect of pruning attention heads instead of neurons. Our search space construction and solution using SA resembles the NeuFair algorithm. NeuFair (Dasu et al., 2024) mitigates unfairness in pre-trained DNNs by identifying a good set of neurons to drop. At each step of the SA search, it creates a copy of the DNN, drops a set of neurons, and analyzes their effect on fairness and model utility. Since it directly performs inference on the unfair DNN, it does not scale to LLMs. The experimental analysis in the paper limits the SA search to small DNNs with at most 384 neurons and low inference times. Even if we apply such an approach for structured pruning at the attention head level instead of unstructured pruning of neurons in LLMs, it would be infeasible, as just one round of inference during the state space exploration would take several minutes. While we quantitatively evaluate against the state-of-the-art attention-pruning strategy for bias reduction, we briefly discuss other bias mitigation strategies in LLMs in three stages of pre-processing, in-processing, and post-processing:
@@
Pre-processing techniques focus on modifying the input data to reduce bias. This includes data curation, where datasets are carefully selected and cleaned to minimize biased content (Bender et al., 2021; Dodge et al., 2021). Another approach is data augmentation, which involves generating additional data to balance the representation of different social groups (Zhou et al., 2023). For instance, counterfactual data generation can be used to create pairs of sentences with perturbed social groups, helping to reduce bias by exposing the model to diverse perspectives (Lu et al., 2020; Panda et al., 2022). Recent works have started using expert models for bias reduction. Orgad and Belinkov (2023) predicted biased samples with an auxiliary model and reweighted them during pre-training. Jeon et al. (2023) employed binary classifiers, or bias experts, to pinpoint biased examples in specific classes.
@@
In-processing methods modify the model’s parameters and optimization process. ADELE (Lauscher et al., 2021) uses adapter layers for parameter-efficient fine-tuning to reduce bias. Regularization techniques can be applied to penalize the model for producing biased outputs (Mozafari et al., 2020). Debiasing word embeddings can help reduce bias in the model’s representation of words by reducing gendered or racial stereotypes (Gonen and Goldberg, 2019; Bolukbasi et al., 2016; Wang et al., 2020). Alternate loss functions such as declustering loss (Garimella et al., 2021) and contrastive loss (Liu et al., 2021; He et al., 2022) have also been shown to be effective for mitigating bias. Zhang et al. (2023) leveraged gradient-based explanations to target sensitive attributes and adjust training for balanced fairness and performance.Dhingra et al. (2023) used style transfer to replace harmful language while maintaining the original meaning.
@@
Post-processing techniques modify the model’s inference behavior after complete training. Gehman et al. (2020) proposed token-blocking methods during decoding to prevent the generation of harmful or biased terms. Hauzenberger et al. (2022) introduced sparse debiasing subnetworks that are trained separately and can be applied to the model at inference time. Qian et al. (2021) performed keyword-based distillation to remove bias during inference. Tokpo and Calders (2022) identify biased tokens and replace them with less stereotypical terms. Post-processing approaches are beneficial since dataset collection and training are expensive processes that we may not be able to repeat if fairness issues are found. Furthermore, post-processing mitigation approaches are sometimes the only viable option with the status quo of utilizing large pre-trained LLMs that take a huge amount of resources to train.
@@ 7. Discussion
Qualitative Examples To better understand how fairness-aware pruning strategies minimize bias, we provide a qualitative example in Table 9. We analyze two prompts belonging to the sex and cisgender subgroups of the Gender group in HolisticBias. We sample 40 tokens from the LLM or until the STOP token is generated, whichever comes first. The responses in all three cases may seem incoherent or incomplete since only 40 tokens are sampled. From the responses, we see that the baseline LLM is biased against the sex subgroup as it generates more toxic content from the prompt compared to the cisgender prompt. FASP and AP do not guarantee zero toxicity. They reduce bias by minimizing the toxicity difference, thereby ensuring similar toxicity across all subgroups. In general, from Table 6, AP is better than other methods at reducing bias across all $33,789$ test set prompts and 8 subgroups in the Gender group of HolisticBias.
@@
Practical Significance of Bias Reduction From Table 6, we see that AP usually has better bias than FASP by $\approx 0.02$ and better perplexity up to $\approx 6$ points or up to $\approx 11\%$. To ensure that this is practically significant, we perform the Cliff’s $\delta$ and Mann-Whitney U statistical tests. We repeat the experiment for 25 random seeds each for AP and FASP. For GPT-J-6B, we observe a $\delta=-0.968$ (95% CI $\approx[-0.994,-0.847]$), indicating a large effect; $\delta$ is negative as lower bias is better. For the Mann-Whitney test, we get $U=10.0$ and (one-sided) $p=2.319\times 10^{-9}$, indicating a 1 in $\approx 400$ million chance that FASP is similar to AP, thus rejecting the null hypothesis that both are from the same distribution. Similarly, for Llama-2-7B, we observe a $\delta=-0.9392$ (95% CI $\approx[-0.984,-0.785]$) which indicates a large effect. For the Mann-Whitney test, we get $U=19$ and (one-sided) $p=6.539\times 10^{-9}$. The perplexity is seed-independent and only depends on the WikiText-2 test split. Apart from achieving lower bias, AP also achieves lower perplexity. State-of-the-art techniques in literature claim a reduction of $\approx 1-2$ points (Sun et al., 2025; Dai et al., 2019) and up to 10% reduction in perplexity (Maini et al., 2024) as significant.
@@
Transferability of Surrogate DNNs While the idea of utilizing surrogate DNNs to speed up the SA search is general, the exact architecture of the DNNs defined in Table 2 is not intended to be transferable to different datasets and fairness metrics. The size and number of layers used in the DNNs depend on the datasets $\mathcal{D}_{\text{bias}}$ and $\mathcal{D}_{\text{ppl}}$ in Algorithm 1 used to train the DNNs. We determine the architecture and hyperparameters using a hyperparameter search to minimize the validation dataset loss. The standard data preprocessing steps outlined in Section 5 can be applied to all datasets to minimize variance. Furthermore, our experiments with the Race attribute (Table 8) demonstrate that the concept of surrogate DNNs extends beyond gender bias.
@@
Limitations We rely on surrogate DNNs to estimate the bias and perplexity. Additionally, we use SA to explore the combinatorial search space. Both contribute to randomness in the search, and our results may be sub-optimal. For example, for GPT-Neo-125M, we can only achieve better perplexity than FASP. We hypothesize this is due to the error in DNN predictions since they are not $100\%$ accurate. Additionally, while SA has statistical guarantees on convergence, it often requires an infeasible number of iterations to reach the optimal state. Another limitation of our approach is that the offline dataset creation to train surrogate DNNs may take time, depending on the LLM and dataset size. For example, the one-time offline data collection of $27,000$ samples and training of surrogate DNNs takes $\approx 1900$ GPU hours in total across multiple GPUs for Llama-2-7B.
@@
Threat to validity To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines by repeating the experiments multiple times and reporting the mean and standard deviation. One challenge is the architecture of DNNs, where we manually performed hyperparameter tuning to find DNNs with ideal accuracy. The results might vary with different architectures, as the SA search relies on them. We also observed outliers in some experiments. For example, in Table 5, we see that the bias of Llama-2-7B is higher when only the bias DNN is used in the cost function compared to the linear combination of both DNNs. This phenomenon was observed after repeating the experiment multiple times. We hypothesize that the relation between bias and perplexity is complex for Llama-2-7B, leaving more analysis to future work. To ensure that our results are generalizable, we performed experiments on six different LLMs of varying sizes. Due to limited computational resources, we only study the HolisticBias and WikiText-2 datasets. Although the state-of-the-art method used a similar experimental setup, it is unknown whether our method will generalize to all LLMs and bias metrics.

### v3

@@ 4. RQ4
Model Baseline AP ($\epsilon=0.5$) FASP (zayed2024fairness) ($\gamma=0.3$) Weight Magnitude (han2016deepcompressioncompressingdeep) Gradient Magnitude (DBLP:journals/corr/abs-1905-10650) Random
@@
Table 6 presents the bias and perplexity of the baseline LLM, AP, and the other strategies. The values in bold correspond to the best results. We see that in 4/6 cases, AP finds states with better bias and perplexity than FASP. The greatest absolute and relative improvement in the bias of $0.18$ and $40.8\%$ is observed for GPT-J-6B. For GPT-Neo-125M, AP’s bias is worse by $\approx 0.02$, but the perplexity is better by $\approx 6$. For Llama-2-7B, AP’s bias is better by $\approx 0.03$ and the perplexity is worse by $\approx 0.3$. Overall, in most cases, AP can reduce bias more while ensuring better perplexity. The other strategies, which do not account for bias, are worse than both AP and FASP [^3] Gradient Magnitude for Llama-2-7B is left blank as its implementation in Hugging Face does not provide an option for computing the gradient of masked heads..
@@
  • The cost function in SA is a linear weighted combination of the two surrogate DNNs. The hyperparameter $\epsilon$ determines how much weight to give to each DNN. Higher values of $\epsilon$ assign more weight to the bias DNN, prioritizing the reduction of bias over the loss of perplexity during the SA search. To visualize the bias-perplexity tradeoff in LLMs, we vary $\epsilon\in[0.3,0.7]$ during the SA search for four LLMs. Figure 2 shows plots of bias versus perplexity of the states for different values of $\epsilon$. States with low bias and high perplexity occur in the upper left region of the graph. States with high bias and low perplexity occur in the lower right region. From the plots, except for outliers that occur due to the randomness of SA, we see that the smaller bubbles that represent lower values of $\epsilon$ occur in the lower right region. Similarly, the larger bubbles moving towards the upper right region represent larger values of $\epsilon$.
@@
  • During surrogate DNN training, increasing $\eta_{bias}$ and $\eta_{ppl}$ uses larger portions of the HolisticBias and WikiText-2 datasets. Larger subsets better represent the true bias/perplexity and can lead to more accurate DNNs. To analyze how accurately the subsets represent the true bias, we sample 25 subsets each for different values of $\eta_{bias}\in[0.1,0.9]$ and plot the error in bias. Figure 3 shows a plot of the error versus $\eta_{bias}$. The dotted line is a regression line fitting all points. Overall, the error decreases as $\eta_{bias}$ increases, confirming that larger validation subsets more accurately estimate the true bias.
@@
  • Increasing the running time (time_limittime\_limit) of the SA search may lead to better states as the search can explore more states. To analyze the effect of running time on the cost and bias/perplexity of best states found, we run the SA search for 5.5 hours with $\epsilon=0.5$ and record the best result at 30-minute intervals. Figure 4 highlights the results for GPT-J-6B. Figure 4(a) shows that the cost decreases over time during the SA search. However, bias/perplexity has a more complicated relationship. From Figure 4(b), we see that bias and perplexity alternate over time, eventually ending with the bias improving more at the expense of perplexity. We observe this because the cost function is a weighted linear function of bias and perplexity. From Figure 2, we see that the best option to control the bias/perplexity trade-off is to vary $\epsilon$.
@@
Answer RQ3: Using larger subsets of the bias/perplexity datasets to train the DNNs can lead to more accurate estimates. Increasing the weightage given to the bias DNN in the cost function reduces bias at the expense of perplexity. Running SA longer may lead to better solutions.
@@
Mitigating one form of bias may unintentionally exacerbate the other biases in the LLM. The HolisticBias metric contains prompts for 13 social biases, and we primarily focus on mitigating gender bias. Similar to zayed2024fairness, to understand how the other social biases are affected by AP, we pick 4 other social biases and evaluate them after repairing gender bias. We pick the best subset of attention heads AP found for gender bias, prune the LLMs, and finally evaluate the repaired LLM on the other biases. Table 7 highlights the effect on other biases before and after repairing gender bias. Note that the perplexity remains identical to the values reported in Table 6 as the perplexity is always evaluated with the WikiText-2 test dataset. We observe that all other social biases decrease for all LLMs. We see a consistent improvement in all cases, even for biases such as age, which are low by default. The highest absolute improvement in bias of 0.29 is observed for GPT-Neo-125M for sexual orientation bias. The highest relative improvement of $65.21\%$ is observed for GPT-J-6B for age bias.
@@ 6. Related Works
The SE community has extensively focused on mitigating unfairness in DNNs, and the literature on mitigating unfairness in LLMs is sparse. DICE (10.1109/ICSE48619.2023.00136) uses a fault localization technique via do logic to identify the top $k$ neurons negatively affecting fairness. FASP uses a similar approach to analyze the effect of pruning attention heads instead of neurons. Our search space construction and solution using SA resembles the NeuFair algorithm. NeuFair (10.1145/3650212.3680380) mitigates unfairness in pre-trained DNNs by identifying a good set of neurons to drop. At each step of the SA search, it creates a copy of the DNN, drops a set of neurons, and analyzes their effect on fairness and model utility. Since it directly performs inference on the unfair DNN, it does not scale to LLMs. The experimental analysis in the paper limits the SA search to small DNNs with at most 384 neurons and low inference times. Even if we apply such an approach for structured pruning at the attention head level instead of unstructured pruning of neurons in LLMs, it would be infeasible, as just one round of inference during the state space exploration would take several minutes. While we quantitatively evaluate against the state-of-the-art attention-pruning strategy for bias reduction, we briefly discuss other bias mitigation strategies in LLMs in three stages of pre-processing, in-processing, and post-processing:
@@
Pre-processing techniques focus on modifying the input data to reduce bias. This includes data curation, where datasets are carefully selected and cleaned to minimize biased content (bender2021dangers; dodge2021documenting). Another approach is data augmentation, which involves generating additional data to balance the representation of different social groups (zhou2023causal). For instance, counterfactual data generation can be used to create pairs of sentences with perturbed social groups, helping to reduce bias by exposing the model to diverse perspectives (lu2020gender; panda2022don). Recent works have started using expert models for bias reduction. orgad-belinkov-2023-blind predicted biased samples with an auxiliary model and reweighted them during pre-training. jeon-etal-2023-improving employed binary classifiers, or bias experts, to pinpoint biased examples in specific classes.
@@
In-processing methods modify the model’s parameters and optimization process. ADELE (lauscher2021sustainable) uses adapter layers for parameter-efficient fine-tuning to reduce bias. Regularization techniques can be applied to penalize the model for producing biased outputs (mozafari2020hate). Debiasing word embeddings can help reduce bias in the model’s representation of words by reducing gendered or racial stereotypes (gonen2019lipstick; bolukbasi2016mancomputerprogrammerwoman; wang2020double). Alternate loss functions such as declustering loss (garimella2021he) and contrastive loss (liu-etal-2021-authors; he-etal-2022-mabel) have also been shown to be effective for mitigating bias. 10.1145/3600211.3604668 leveraged gradient-based explanations to target protected attributes and adjust training for balanced fairness and performance. dhingra2023queer used style transfer to replace harmful language while maintaining the original meaning.
@@
Post-processing techniques modify the model’s inference behavior after complete training. gehman2020realtoxicityprompts proposed token-blocking methods during decoding to prevent the generation of harmful or biased terms. hauzenberger2022modular introduced sparse debiasing subnetworks that are trained separately and can be applied to the model at inference time. qian2021counterfactual performed keyword-based distillation to remove bias during inference. tokpo2022text identify biased tokens and replace them with less stereotypical terms. Post-processing approaches are beneficial since dataset collection and training are expensive processes that we may not be able to repeat if fairness issues are found. Furthermore, post-processing mitigation approaches are sometimes the only viable option with the status quo of utilizing large pre-trained LLMs that take a huge amount of resources to train.
@@ 7. Discussion
Qualitative Examples. To better understand how fairness-aware pruning strategies minimize bias, we provide a qualitative example in Table 9. We analyze two prompts belonging to the sex and cisgender subgroups of the Gender group in HolisticBias. We sample 40 tokens from the LLM or until the STOP token is generated, whichever comes first. The responses in all three cases may seem incoherent or incomplete since only 40 tokens are sampled. From the responses, we see that the baseline LLM is biased against the sex subgroup as it generates more toxic content from the prompt compared to the cisgender prompt. FASP and AP do not guarantee zero toxicity. They reduce bias by minimizing the toxicity difference, thereby ensuring similar toxicity across all subgroups. In general, from Table 6, AP is better than other methods at reducing bias across all $33,789$ test set prompts and 8 subgroups in the Gender group of HolisticBias.
@@
Practical Significance of Bias Reduction. From Table 6, we see that AP usually has better bias than FASP by $\approx 0.02$ and better perplexity up to $\approx 6$ points or up to $\approx 11\%$. To ensure that this is practically significant, we perform the Cliff’s $\delta$ and Mann-Whitney U statistical tests. We repeat the experiment for 25 random seeds each for AP and FASP. For GPT-J-6B, we observe a $\delta=-0.968$ (95% CI $\approx[-0.994,-0.847]$), indicating a large effect; $\delta$ is negative as lower bias is better. For the Mann-Whitney test, we get $U=10.0$ and (one-sided) $p=2.319\times 10^{-9}$, indicating a 1 in $\approx 400$ million chance that FASP is similar to AP, thus rejecting the null hypothesis that both are from the same distribution. Similarly, for Llama-2-7B, we observe a $\delta=-0.9392$ (95% CI $\approx[-0.984,-0.785]$) which indicates a large effect. For the Mann-Whitney test, we get $U=19$ and (one-sided) $p=6.539\times 10^{-9}$. The perplexity is seed-independent and only depends on the WikiText-2 test split. Apart from achieving lower bias, AP also achieves lower perplexity. State-of-the-art techniques in literature claim a reduction of $\approx 1-2$ points (sun2025cursedepthlargelanguage; dai2019transformerxlattentivelanguagemodels) and up to 10% reduction in perplexity (maini2024rephrasingwebrecipecompute) as significant.
@@
Transferability of Surrogate DNNs. While the idea of utilizing surrogate DNNs to speed up the SA search is general, the exact architecture of the DNNs defined in Table 2 is not intended to be transferable to different datasets and fairness metrics. The size and number of layers used in the DNNs depend on the datasets $\mathcal{D}_{\text{bias}}$ and $\mathcal{D}_{\text{ppl}}$ in Algorithm 1 used to train the DNNs. We determine the architecture and hyperparameters using a hyperparameter search to minimize the validation dataset loss. The standard data preprocessing steps outlined in Section 5 can be applied to all datasets to minimize variance. Furthermore, our experiments with the Race attribute (Table 8) demonstrate that the concept of surrogate DNNs extends beyond gender bias.
@@
Limitations. We rely on surrogate DNNs to estimate the bias and perplexity. Additionally, we use SA to explore the combinatorial search space. Both contribute to randomness in the search, and our results may be sub-optimal. For example, for GPT-Neo-125M, we can only achieve better perplexity than FASP. We hypothesize this is due to the error in DNN predictions since they are not $100\%$ accurate. Additionally, while SA has statistical guarantees on convergence, it often requires an infeasible number of iterations to reach the optimal state. Another limitation of our approach is that the offline dataset creation to train surrogate DNNs may take time, depending on the LLM and dataset size. For example, the one-time offline data collection of $27,000$ samples and training of surrogate DNNs takes $\approx 1900$ GPU hours in total across multiple GPUs for Llama-2-7B.
@@
Threats to validity. To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines by repeating the experiments multiple times and reporting the mean and standard deviation. One challenge is the architecture of DNNs, where we manually performed hyperparameter tuning to find DNNs with ideal accuracy. The results might vary with different architectures, as the SA search relies on them. We also observed outliers in some experiments. For example, in Table 5, we see that the bias of Llama-2-7B is higher when only the bias DNN is used in the cost function compared to the linear combination of both DNNs. This phenomenon was observed after repeating the experiment multiple times. We hypothesize that the relationship between bias and perplexity is complex for Llama-2-7B, leaving more analysis to future work. To ensure that our results are generalizable, we performed experiments on six different LLMs of varying sizes. Due to limited computational resources, we only study the HolisticBias and WikiText-2 datasets. Although the state-of-the-art method used a similar experimental setup, it is unknown whether our method will generalize to all LLMs and bias metrics.